OpenStudy (anonymous):

(Pre-calculus) In 1987, the number of AIDS cases in the US was estimated to be 50,000. In 2005, this number was estimated to be 1,000,000. Assume the trend continued. [leave answers in terms of e and ln, etc] a) Set up a function of the form N(t) = N0 x e^kt b) According to your model, how many cases occurred in 2000? c) When, according to your model, will there be 5,000,000 cases?

OpenStudy (dumbcow):

for part a) you need to solve for k assign values for N_0 (initial number) and N(t) where t = 18 (2005 - 1987 = 18)

OpenStudy (dumbcow):

\[\rightarrow 1,000,000 = 50,000*e^{18k}\]

OpenStudy (anonymous):

Ok, great thanks!! And then my constant should have an ln answer in it right?

OpenStudy (dumbcow):

correct... k = ln(20) / 18

OpenStudy (anonymous):

ok great!! and then for part c, I would put: 5,000,000No =50,000ekt?

OpenStudy (dumbcow):

almost, N_0 = 50,000 so you don't need it on left hand side then solve for t

OpenStudy (anonymous):

Ok, got it! So get rid of the No and plug in 5,00,000 = 50,000e(ln20/18)t . e cancels out ln, so it would be 5mil=50000(20)(t/18)!

OpenStudy (dumbcow):

yep except the (t/18) remains an exponent...do Not multiply \[e^{\ln 20 / 18 t} = (e^{\ln 20})^{t/18} = 20^{t/18}\]

OpenStudy (anonymous):

Oh..don't multiply....could I possibly do log base 20 to both sides to get solve for t?

OpenStudy (dumbcow):

you could...i think its less confusing to stick with ln , then use log property to bring exponent out front

OpenStudy (anonymous):

Ah okay! got it!! Thanks for all the much needed help. I'm taking a summer pre-calculus college class right now and our final is next week!! You're awesome!

OpenStudy (dumbcow):

your welcome..good luck

OpenStudy (anonymous):

thanks!! I don't want to call you dumbcow because you're definitely way better than that!